OpenStudy (sloppycanada):

Find the period of f(x) = sin(1/2)x

OpenStudy (sloppycanada):

\[f(x)= \sin \frac{ 1 }{ 2 }x\] Isn't this just \[\frac{ 1 }{ 2 }\]

OpenStudy (sloppycanada):



OpenStudy (baru):

its \(\frac{2\pi}{1/2}\)

OpenStudy (baru):

you understand how to find time period?

OpenStudy (sloppycanada):

I do, but \[ \frac{ 2\pi }{ 1/2 }\] isn't an option.

OpenStudy (baru):

is \(4\pi\) an option?

OpenStudy (baru):

@sloppycanada

OpenStudy (baru):

\( 2\pi \div 1/2 =2\pi \times 2\)

OpenStudy (sloppycanada):

Yes @baru Sorry, someone was at the door.

OpenStudy (sloppycanada):

What is the period of the function f(x)= cos2x? This is just \[\frac{ 1 }{ 2 }\pi\] Correct?

OpenStudy (baru):

its \(2\pi/2\)

OpenStudy (sloppycanada):

Again, not an option... So, 4 again?

OpenStudy (anonymous):

\(2 \pi/2=\pi\)

OpenStudy (baru):

if ur function is \[Asin(\omega x)\] time period \[T=\frac{2\pi }{\omega}\]